OpenStudy (anonymous):

Each of the letters in the word MATHEMATICS is represented by a letter tile in a bag. Stacey picks three letters, without replacement. To the nearest whole percent, which is the probability that she'll get all vowels? answer choices: 2% -3%-5%-6%

OpenStudy (anonymous):

you find the percentage of picking a vowel so 4/11 then you find the percentage of picking 3 out of the 4 so 3/4 then multiply them

OpenStudy (anonymous):

you there?

OpenStudy (kropot72):

The probability of picking a vowel on the first selection is 4/11, on the second selection is 3/10 and on the third selection is 2/9. the percentage probability of picking all vowels is: \[\frac{4}{11}\times \frac{3}{10}\times \frac{2}{9}\times \frac{100}{1}=?\]